Circular No. 8621

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 T5 (BROUGHTON) John Broughton, Reedy Creek, Qld., Australia, reports his discovery

of a comet on CCD images taken with a 0.51-m f/2.7 reflector (discovery observation below). Other observers have confirmed the object’s cometary appearance from their own CCD images, including J. Young (Table Mountain, 0.6-m reflector, co-added exposures taken on Oct. 24.2 UT; prominent, round 8′′ coma with a 16′′ tail spanning p.a. 20^{o}–30^{o}), R. H. McNaught (Siding Spring, 1.0-m reflector, 60-s exposures taken on Oct. 24.45; very strongly condensed condensation of size 2′′.5 FWHM with a faint 12′′ tail in p.a. 40^{o}), and J. E. McGaha (Tucson, 0.62-m reflector, five 180-s co-added frames taken on Oct. 25.2; round, bright inner coma of diameter ∼ 7′′ and a very faint outer coma extending to ∼ 20′′, no tail).

2005 UT  α_{2000}  δ_{2000}  Mag.
Oct. 9.57058  0^{h}06^{m}38.^{s}05  −27^{o}22′54′′.2  18.5

The available astrometry (prediscovery observations back to Aug. 26), the following orbital elements, and an ephemeris appear on MPEC 2005-U53.

T = 2005 Nov. 3.3917 TT  ω = 304.^{o}4863
e = 0.552357  Ω = 57.0701 }2000.0
q = 3.247088 AU  i = 21.3730
a = 7.253744 AU n^{o} = 0.0504499 P = 19.536 years

SUPERNOVAE 2005hc, 2005hd, 2005he Three supernovae have been reported: 2005hc in MCG +00-6-3 by

J. Frieman (Fermilab and University of Chicago), on behalf of the Sloan Digitial Sky Survey II collaboration; 2005hd by the “Nearby Supernova Factory” collaboration; and 2005he in MCG +06-49-68 by D. R. Madison and W. Li on unfiltered LOSS/KAIT images (cf. IAUC 8616).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005hc Oct. 12  1^{h}56^{m}47.^{s}94 − 0^{o}12′49′′.2 18.5 8′′.4 W, 3′′.9 S
2005hd Oct. 21.5 4 16 12.37 −13 13 11.6 17.8 –
2005he Oct. 24.23 22 39 45.44 +34 22 43.0 18.6 1′′.3 W, 5′′.5 S

SN 2005hc is a type-Ia supernova (details on CBETs 259 and 260). SN 2005hd is a type-II supernova (details on CBET 263). Additional KAIT magnitudes for 2005he: Sept. 29.24 UT, [19.5; Oct. 16.22, [18.5; 25.22, 18.7.

2005 October 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green